OpenStudy (anonymous):

11sqrt{20} + 4sqrt{45}

OpenStudy (anonymous):

\[11\sqrt{20}+4\sqrt{45}\]

OpenStudy (anonymous):

explain steps please

OpenStudy (anonymous):

ok so you have the problem \[11\sqrt{20}+4\sqrt{45}\] you break down the sqrts like so \[22\sqrt{5}+12\sqrt{5}\] then add them into one sqrt like so \[34\sqrt{5}\] which is equal about 76.02 check work \[11\sqrt{20}+4\sqrt{45}=76.02\] and \[34\sqrt{5}=76.02\]

OpenStudy (anonymous):

so i have to square each number?

OpenStudy (anonymous):

do you under stand how i broke each one down?

OpenStudy (anonymous):

then I add. ok this where i got lost on the decimal stuff. is that like the three decimal places thing? I understood all up to the decimal thing

OpenStudy (anonymous):

the 76.02?

OpenStudy (anonymous):

yes how did you get that? i rem doing some work with the square roots that said simplify to three decimal places

OpenStudy (anonymous):

yea like what is the value to 3 decimal places. it said i could use a calculator but i couldnt get it on the calculator

OpenStudy (anonymous):

ok so all i did was simply it on paper then used a calculator to find out the 76.02 its accually like 76.02631123

OpenStudy (anonymous):

so 76.026

OpenStudy (anonymous):

what do i push in on the calculator to find that i did the sqrt symbol then number but it doesnt do anything

OpenStudy (anonymous):

uhhh what calculator do you have?

OpenStudy (anonymous):

fx-300es

OpenStudy (anonymous):

casio

OpenStudy (anonymous):

it should just jsut be like 35*sqrt(5) or (11*sqrt(20))+(4*sqrt(45))

OpenStudy (anonymous):

sorry if i cant help more than that i use a ti-84 plus so im not 100% sure what buttons are where on yours

OpenStudy (anonymous):

yea its hard to tell when there are different calculators. thanks though

OpenStudy (anonymous):

sorry wish i could be more help

OpenStudy (anonymous):

oh no your helping just fine. id rather have it explained to me than just give me the answers

OpenStudy (anonymous):

do you need me to explain something else?

OpenStudy (anonymous):

im working on one right now its \[\sqrt[4]x \times \sqrt[5]7x{}\]

OpenStudy (anonymous):

do i add 4 and 5

OpenStudy (anonymous):

hmmm.....give me a min to think

OpenStudy (anonymous):

http://www.brightstorm.com/math/algebra-2/roots-and-radicals/multiplying-radicals-of-different-roots/ watch this video it explains how to do that better than i can

OpenStudy (anonymous):

hope that helps some

OpenStudy (anonymous):

\[\sqrt[20]7x{?}\]

OpenStudy (anonymous):

it helps, but when its not in the same form as my problem i dont knwo how to do it.